# Offline-online approximation of multiscale eigenvalue problems with random defects

### v1

@@ 4.1 Preliminary estimates for the MLOD
$\begin{split}[a(v^{\mathrm{ms}}_{H,\mathfrak{k}},v_{H,\mathfrak{k}})-\lambda]\leq\norm{u-v^{\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}-\lambda\norm{u-v^{\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{L^{2}(\Omega)}+\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{H,\mathfrak{k}}}_{\mathcal{A}}\norm{v_{H,\mathfrak{k}}}_{\mathcal{A}}.\end{split}$
Using this inequality and our definitions, we obtain that
$\begin{split}\frac{\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}{\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}-\lambda&=[\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})-\lambda]+\big{[}\frac{\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}{\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}-\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})\big{]}\\ &=[\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})-\lambda]+\big{[}\frac{\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}{\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}\norm{v^{\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{L^{2}(\Omega)}-\frac{\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}{\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})\big{]}\\ &=[\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})-\lambda]+\frac{\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}{\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}\big{[}{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})-\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})\big{]}\\ &=a(v^{\mathrm{ms}}_{H,\mathfrak{k}},v_{H,\mathfrak{k}})-\lambda+\frac{\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}{\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}\big{[}{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})-\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})\big{]}\\ &\leq\norm{u-v^{\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}-\lambda\norm{u-v^{\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{L^{2}(\Omega)}+\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{H,\mathfrak{k}}}_{\mathcal{A}}\norm{v_{H,\mathfrak{k}}}_{\mathcal{A}}\\ &\hskip 56.9055pt+\frac{\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}{\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}[{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})-\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})].\qed\end{split}$
@@ 4.2 Main results for the MLOD method
Let $\lambda^{(l)}$ be an eigenvalue of multiplicity $r$, i.e., $\lambda^{(l)}=\cdots=\lambda^{(l+r-1)}$ with the corresponding eigenspace $E_{l+j},\;{j=0,1,\cdots,r-1}$ spanned by the $r$-many orthonormal basis functions with $\norm{u^{(l+j)}}_{\mathcal{H}^{1}}=1$. Suppose the pairs $({\lambda}^{(l)}_{H,\mathfrak{k}},{u}_{H,\mathfrak{k}}^{(l),\mathrm{ms}}),\cdots,({\lambda}^{(l+r-1)}_{H,\mathfrak{k}},{u}_{H,\mathfrak{k}}^{(l+r-1),\mathrm{ms}})$ approximate the LOD problem in (7) with $\norm{{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}}_{L^{2}(\Omega)}=1$ and the pairs $(\tilde{\lambda}^{(l)}_{H,\mathfrak{k}},\tilde{u}_{H,\mathfrak{k}}^{(l),\mathrm{ms}}),\cdots,(\tilde{\lambda}^{(l+r-1)}_{H,\mathfrak{k}},\tilde{u}_{H,\mathfrak{k}}^{(l+r-1),\mathrm{ms}})$ approximate the problem in (9) with $\norm{\tilde{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}}_{L^{2}(\Omega)}=1\;\forall j$. Further assume $\overline{E}^{\mathfrak{k}}_{l+j}$ and $\widetilde{E}^{\mathfrak{k}}_{l+j}$ to be the eigenspace spanned by the $r$-many basis functions corresponding to ${\lambda}^{(l+j)}_{H,\mathfrak{k}}$ and $\tilde{\lambda}^{(l+j)}_{H,\mathfrak{k}},\;\forall j$ respectively. If $l+r-1\leq N$, then, there exists a function $u^{(l+j)}\in E_{l+j}$ such that, for all $j=0,1,\cdots,r-1$,
$\norm{u^{(l+j)}-\tilde{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}}_{\mathcal{A}}\lesssim H^{2}+g(\mathfrak{k}).$
@@
For any $j=0,1,\cdots,r-1$, with $u^{(l+j)}\in E_{l+j}$ we have that
@@
$\begin{split}\norm{u^{(l+j)}-\tilde{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}}_{\mathcal{A}}&\leq\norm{u^{(l+j)}-{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}}_{\mathcal{A}}+\norm{{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}-\tilde{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}}_{\mathcal{A}}\\ &\leq\norm{u^{(l+j)}-{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}}_{\mathcal{A}}+\norm{{u}_{H,\mathfrak{k}}^{(l+j)}-\tilde{u}_{H,\mathfrak{k}}^{(l+j)}}_{\mathcal{A}}+\norm{\mathscr{C}_{\mathfrak{k}}({u}_{H,\mathfrak{k}}^{(l+j)}-\tilde{u}_{H,\mathfrak{k}}^{(l+j)})}_{\mathcal{A}}\\ &\leq\norm{u^{(l+j)}-{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}}_{\mathcal{A}}+2\norm{{u}_{H,\mathfrak{k}}^{(l+j)}-\tilde{u}_{H,\mathfrak{k}}^{(l+j)}}_{\mathcal{A}}+\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})(\overline{u}_{H,\mathfrak{k}}^{(l+j)}-\tilde{u}_{H,\mathfrak{k}}^{(l+j)})}_{\mathcal{A}}\\ &\lesssim H^{2}+\Big{(}g(\mathfrak{k})\Big{)}^{1/2}H+(2+g(\mathfrak{k}))\norm{{u}_{H,\mathfrak{k}}^{(l+j)}-\tilde{u}_{H,\mathfrak{k}}^{(l+j)}}_{\mathcal{A}}\end{split}$
where the second last step follows from the stability of $\mathscr{C}$ in the energy norm. The last step is due to (8) and Theorem 2.2. The second term can be estimated using the compact operators. Suppose in Lemma 3.1, $M1$ is given by the LOD solving (7) and $M2$ is given by the MLOD solving (9). Then, with $a_{1}={a}_{\mathfrak{k}}$, $a_{2}=\widetilde{a}_{\mathfrak{k}}$, $m_{1}={m}_{\mathfrak{k}}$, $m_{2}=\widetilde{m}_{\mathfrak{k}}$, $T^{1}_{H}={T}^{\mathfrak{k}}_{H}$ and $T^{2}_{H}=\widetilde{T}^{\mathfrak{k}}_{H}$, it yields
@@
$\begin{split}\norm{{T}^{\mathfrak{k}}_{H}f^{(l+j)}-\widetilde{T}^{\mathfrak{k}}_{H}f^{(l+j)}}^{2}_{\mathcal{A}}&\lesssim({m}_{\mathfrak{k}}-\widetilde{m}_{\mathfrak{k}})(f^{(l+j)},{T}^{\mathfrak{k}}f^{(l+j)}-\widetilde{T}^{\mathfrak{k}}_{H}f^{(l+j)})\\ &\hskip 59.75095pt+(\widetilde{a}_{\mathfrak{k}}-{a}_{\mathfrak{k}})(\widetilde{T}^{\mathfrak{k}}f^{(l+j)},{T}^{\mathfrak{k}}_{H}f^{(l+j)}-\widetilde{T}^{\mathfrak{k}}_{H}f^{(l+j)}).\end{split}$
Since $\widetilde{T}^{\mathfrak{k}}_{H}f^{l+j}$ solves for (7), by Lemma 4.1 and Lemma 4.2, we obtain that
$\begin{split}\norm{{T}^{\mathfrak{k}}_{H}f^{(l+j)}-\widetilde{T}^{\mathfrak{k}}_{H}f^{(l+j)}}_{\mathcal{A}}&\lesssim\norm{f}_{\mathcal{A}}H^{2}+\norm{\widetilde{T}^{\mathfrak{k}}_{H}f^{(l+j)}}_{\mathcal{A}}g(\mathfrak{k}).\end{split}$
@@
Since ${T}^{\mathfrak{k}}_{H}f^{(l+j)}$, $\widetilde{T}^{\mathfrak{k}}_{H}f^{(l+j)}\in V_{H}$, there exist some ${v}_{0}$, $\tilde{v}_{0}\in V_{H}$ such that ${T}^{\mathfrak{k}}_{H}f^{(l+j)}={v}_{0}$ and $\widetilde{T}^{\mathfrak{k}}_{H}f^{(l+j)}=\tilde{v}_{0}$ solving (20) and (21) for LOD and MLOD respectively. On the other hand, (7) and (9) solve for eigensolutions on $V_{H}$ for the LOD and MLOD respectively. Moreover, the pair (20) and (7) (similarly (21) and (9)) produce the same eigenfunctions and reciprocal eigenvalues to the LOD and MLOD problems. Therefore, for $f\in E_{(l+j)}$ with $\norm{f^{(l+j)}}_{\mathcal{H}^{1}(\Omega)}=1$, by the analysis of [BO90] we obtain that
@@
$\begin{split}\norm{{u}^{(l+j)}_{H,\mathfrak{k}}-\tilde{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}&\lesssim H^{2}+g(\mathfrak{k})\norm{\tilde{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}.\end{split}$ (25)
@@
We are left to bound $\norm{\tilde{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}$. For $\mathfrak{k}$ large enough, by the coercivity of $\widetilde{a}(\cdot,\cdot)$ [EGH15], for any $v\in\mathcal{H}^{1}_{\#}$
@@
$\begin{split}\norm{\tilde{u}_{H,\mathfrak{k}}^{(l+j)}}_{\mathcal{A}}^{2}\leq\;\absolutevalue{\widetilde{a}_{\mathfrak{k}}(\tilde{u}_{H,\mathfrak{k}}^{(l+j)},\tilde{u}_{H,\mathfrak{k}}^{(l+j)})}=&\;\absolutevalue{\tilde{\lambda}_{H,\mathfrak{k}}^{l}\widetilde{m}_{\mathfrak{k}}(\tilde{u}_{H,\mathfrak{k}}^{(l+j)},\tilde{u}_{H,\mathfrak{k}}^{(l+j)})}\leq\widetilde{C}_{\lambda}\absolutevalue{\lambda^{(l)}}\absolutevalue{m(\tilde{u}_{H,\mathfrak{k}}^{(l+j)},\tilde{u}_{H,\mathfrak{k}}^{(l+j)})}=\widetilde{C}_{\lambda}\absolutevalue{\lambda^{(l)}}\norm{\tilde{u}_{H,\mathfrak{k}}^{(l+j)}}^{2}_{L^{2}(\Omega)}.\end{split}$ (26)
The third step follows due to |$\absolutevalue{\tilde{\lambda}^{(l)}_{H,\mathfrak{k}}}\leq c\absolutevalue{\lambda^{(l)}},$ for some $c\in\mathbb{R}$ by Babuška-Osborn theory because Lemma 3.1 and the known analysis of the source problem (see above) in particular yield convergence of $\tilde{\lambda}^{(l)}_{H,\mathfrak{k}}$ to $\lambda^{(l)}$ for $H\to 0$ and $\mathfrak{k}\to\infty$. Since $\norm{\tilde{u}_{H,\mathfrak{k}}^{(l+j)}}_{L^{2}(\Omega)}=\norm{I_{H}\tilde{u}_{H,\mathfrak{k}}^{(l+j),ms}}_{L^{2}(\Omega)}\leq\widetilde{C}_{\mathfrak{k}}\norm{\tilde{u}_{H,\mathfrak{k}}^{(l+j),ms}}_{L^{2}(\Omega)}=\widetilde{C}_{\mathfrak{k}}$ by definition, $\norm{\tilde{u}_{H,\mathfrak{k}}^{(l+j)}}_{\mathcal{A}}$ is clearly bounded above by a positive constant. Therefore, it holds that
$\begin{split}\norm{u^{(l+j)}-\tilde{u}_{H,\mathfrak{k}}^{(l+j),ms}}_{\mathcal{A}}&\lesssim 2H^{2}+2g(\mathfrak{k})+H^{2}g(\mathfrak{k})+g^{2}(\mathfrak{k})+\Big{(}g(\mathfrak{k})\Big{)}^{1/2}H+H^{2}.\end{split}$

### v2

@@ 4.1 Preliminary estimates for the MLOD
$\begin{split}a\left(v^{\mathrm{ms}}_{H,\mathfrak{k}},v_{H,\mathfrak{k}}\right)-\lambda\leq\norm{u-v^{\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}-\lambda\norm{u-v^{\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{L^{2}(\Omega)}+\norm{\left(\mathscr{C}-\mathscr{C}_{\mathfrak{k}}\right)v_{H,\mathfrak{k}}}_{\mathcal{A}}\norm{v_{H,\mathfrak{k}}}_{\mathcal{A}}.\end{split}$
Recall that $\widetilde{a}_{\mathfrak{k}}(\cdot,\cdot)=a\left((\cdot)^{\mathrm{ms}}_{\mathfrak{k}},(\cdot)\right)$ and $\widetilde{m}_{\mathfrak{k}}(\cdot,\cdot)=m((\cdot),(\cdot))=\norm{(\cdot)}^{2}_{L^{2}(\Omega)}$. Then, together with the above inequality, we obtain that
$\begin{split}\frac{\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}{\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}-\lambda&=[\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})-\lambda]+\left[\frac{\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}{\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}-\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})\right]\\ &=[\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})-\lambda]+\left[\frac{\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}{\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}\norm{v^{\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{L^{2}(\Omega)}-\frac{\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}{\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})\right]\\ &=[\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})-\lambda]+\frac{\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}{\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}\big{[}{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})-\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})\big{]}\\ &=a(v^{\mathrm{ms}}_{H,\mathfrak{k}},v_{H,\mathfrak{k}})-\lambda+\frac{\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}{\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}\big{[}{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})-\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})\big{]}\\ &\leq\norm{u-v^{\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}-\lambda\norm{u-v^{\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{L^{2}(\Omega)}+\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})v_{H,\mathfrak{k}}}_{\mathcal{A}}\norm{v_{H,\mathfrak{k}}}_{\mathcal{A}}\\ &\hskip 56.9055pt+\frac{\widetilde{a}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}{\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})}\left[{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})-\widetilde{m}_{\mathfrak{k}}(v_{H,\mathfrak{k}},v_{H,\mathfrak{k}})\right].\qed\end{split}$
@@ 4.2 Main results for the MLOD method
Let $\lambda^{(l)}$ be the $l$-th eigenvalue of the continuous problem (3) with multiplicity $r$, i.e., $\lambda^{(l)}=\lambda^{(l+1)}=\cdots=\lambda^{(l+r-1)}$, and let the corresponding eigenspace $E^{(l+j)}\subset\mathcal{H}^{1}_{\#}(\Omega),$ for ${j=0,1,\cdots,r-1}$ be spanned by the $r$-many orthonormal basis functions. Let $\left\{\tilde{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}\right\}_{j=0}^{r-1}\subset V^{\mathrm{ms}}_{H,\mathfrak{k}}$ be an orthonormal set of multiscale eigenfunctions approximating the MLOD problem in (8) with $\norm{\tilde{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}}_{L^{2}(\Omega)}=1\;\text{ for every }j$. Assume $l+r-1\leq N$, where $N=\mathrm{dim}(V_{H})$. Then, for each $j=0,1,\cdots,r-1$, there exists a function $\tilde{u}^{(l+j)}\in E^{(l+j)}$ with $\norm{\tilde{u}^{(l+j)}}_{L^{2}(\Omega)}=1$ such that
$\norm{\tilde{u}^{(l+j)}-\tilde{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}}_{\mathcal{A}}\lesssim H^{2}+g(\mathfrak{k}).$
The hidden constant depends only on the spectral bounds, domain $\Omega$, multiplicity $r$ and the corresponding continuous eigenvalue $\lambda^{(l)}$.
@@
We further note that the choice of $\tilde{u}^{(l+j)}\in E^{(l+j)}$ such that $\norm{\tilde{u}^{(l+j)}-\tilde{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}}_{\mathcal{A}}\to 0$ as $H\to 0$ and $\mathfrak{k}\to\infty$ is made a posteriori, i.e., for each computed eigenfunction, a corresponding function in the exact eigenspace is chosen to minimize the energy norm error. Such matching is guaranteed by classical results for perturbations of compact operators as discussed in [Cha83, BO91]. This standard approach allows us to make individual per-function estimates even in the presence of eigenvalue multiplicities, since the continuous and discrete eigenspaces converge as subspaces under mesh refinement and localization.
@@
For any $j=0,1,\cdots,r-1$, with $\tilde{u}^{(l+j)}\in E^{(l+j)}$ we can write
$\norm{\tilde{u}^{(l+j)}-\tilde{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}}_{\mathcal{A}}\leq\norm{\tilde{u}^{(l+j)}-{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}}_{\mathcal{A}}+\norm{{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}-\tilde{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}}_{\mathcal{A}}$
with ${u}_{H,\mathfrak{k}}^{(l+j)}\in V^{\text{ms}}_{H}$ being the approximate solution to the LOD problem in (6). Since $V^{\text{ms}}_{H,\mathfrak{k}}\in\mathcal{H}^{1}_{\#}$ and $\tilde{u}^{(l+j)}\in\mathcal{H}^{1}_{\#}$, all terms can be interpreted within the same space. We can then proceed as follows using the triangle inequality
@@
$\begin{split}\norm{\tilde{u}^{(l+j)}-\tilde{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}}_{\mathcal{A}}&\leq\norm{\tilde{u}^{(l+j)}-{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}}_{\mathcal{A}}+\norm{{u}_{H,\mathfrak{k}}^{(l+j)}-\tilde{u}_{H,\mathfrak{k}}^{(l+j)}}_{\mathcal{A}}+\norm{\mathscr{C}_{\mathfrak{k}}\left({u}_{H,\mathfrak{k}}^{(l+j)}-\tilde{u}_{H,\mathfrak{k}}^{(l+j)}\right)}_{\mathcal{A}}\\ &\leq\norm{\tilde{u}^{(l+j)}-{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}}_{\mathcal{A}}+2\norm{{u}_{H,\mathfrak{k}}^{(l+j)}-\tilde{u}_{H,\mathfrak{k}}^{(l+j)}}_{\mathcal{A}}+\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})\left({u}_{H,\mathfrak{k}}^{(l+j)}-\tilde{u}_{H,\mathfrak{k}}^{(l+j)}\right)}_{\mathcal{A}}\\ &\lesssim H^{2}+\left(g(\mathfrak{k})\right)^{1/2}H+\left(2+g(\mathfrak{k})\right)\norm{{u}_{H,\mathfrak{k}}^{(l+j)}-\tilde{u}_{H,\mathfrak{k}}^{(l+j)}}_{\mathcal{A}},\end{split}$
where the second last step follows from the stability of $\mathscr{C}$ in the energy norm. In the last step we used the estimates from (7) for the $l$-th eigenvalue with multiplicity $r$. These estimates for $r>1$ are justified by the analysis in [MP15]. Furthermore, the function $\tilde{u}^{(l+j)}$ is chosen to minimize the energy norm error with respect to the MLOD eigenfunction $\tilde{u}^{(l+j),\text{ms}}_{H,\mathfrak{k}}$. Therefore, it is not necessarily the best matching function for the LOD approximation ${u}^{(l+j),\text{ms}}_{H,\mathfrak{k}}$. However, since all approximate eigenspaces converge to $E^{(l+j)}$ in norm, the estimates used in (7) are valid for estimating the first term in the last step above, and the convergence rate is preserved up to a constant. We have further used Theorem 2.2 to approximate $\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})\left({u}_{H,\mathfrak{k}}^{(l+j)}-\tilde{u}_{H,\mathfrak{k}}^{(l+j)}\right)}_{\mathcal{A}}$ to arrive at the last step. What is left to bound is $\norm{{u}_{H,\mathfrak{k}}^{(l+j)}-\tilde{u}_{H,\mathfrak{k}}^{(l+j)}}_{\mathcal{A}}$, which we can estimate by using compact operators. Suppose in Lemma 3.1, $M1$ is given by the LOD solving (6) and $M2$ is given by the MLOD solving (8). Then, with $a_{1}={a}_{\mathfrak{k}}$, $a_{2}=\widetilde{a}_{\mathfrak{k}}$, $m_{1}={m}_{\mathfrak{k}}$, $m_{2}=\widetilde{m}_{\mathfrak{k}}$, $T^{1}_{H}={T}^{\mathfrak{k}}_{H}$ and $T^{2}_{H}=\widetilde{T}^{\mathfrak{k}}_{H}$, for the $(l+j)$-th eigenfunction it yields
@@
$\begin{split}\norm{{T}^{\mathfrak{k}}_{H}f^{(l+j)}-\widetilde{T}^{\mathfrak{k}}_{H}f^{(l+j)}}^{2}_{\mathcal{A}}&\lesssim({m}_{\mathfrak{k}}-\widetilde{m}_{\mathfrak{k}})(f^{(l+j)},{T}^{\mathfrak{k}}_{H}f^{(l+j)}-\widetilde{T}^{\mathfrak{k}}_{H}f^{(l+j)})\\ &\hskip 59.75095pt+(\widetilde{a}_{\mathfrak{k}}-{a}_{\mathfrak{k}})(\widetilde{T}^{\mathfrak{k}}_{H}f^{(l+j)},{T}^{\mathfrak{k}}_{H}f^{(l+j)}-\widetilde{T}^{\mathfrak{k}}_{H}f^{(l+j)}).\end{split}$
Since $\widetilde{T}^{\mathfrak{k}}_{H}f^{(l+j)}$ solves for (6), by Lemma 4.1 and Lemma 4.2, we obtain that
$\begin{split}\norm{{T}^{\mathfrak{k}}_{H}f^{(l+j)}-\widetilde{T}^{\mathfrak{k}}_{H}f^{(l+j)}}_{\mathcal{A}}&\lesssim\norm{f^{(l+j)}}_{\mathcal{A}}H^{2}+\norm{\widetilde{T}^{\mathfrak{k}}_{H}f^{(l+j)}}_{\mathcal{A}}g(\mathfrak{k}).\end{split}$
@@
Since ${T}^{\mathfrak{k}}_{H}f^{(l+j)}$, $\widetilde{T}^{\mathfrak{k}}_{H}f^{(l+j)}\in V_{H}$, there exist some ${v}_{0}$, $\tilde{v}_{0}\in V_{H}$ such that ${T}^{\mathfrak{k}}_{H}f^{(l+j)}={v}_{0}$ and $\widetilde{T}^{\mathfrak{k}}_{H}f^{(l+j)}=\tilde{v}_{0}$ solving (17) and (18) for LOD and MLOD respectively. On the other hand, (6) and (8) solve for eigensolutions on $V_{H}$ for the LOD and MLOD respectively. Moreover, the pair (17) and (6) (similarly (18) and (8)) produce the same eigenfunctions and reciprocal eigenvalues to the LOD and MLOD problems. Therefore, for $f^{(l+j)}\in E^{(l+j)}$ with $\norm{f^{(l+j)}}_{L^{2}(\Omega)}=1$, the analysis of compact operators for approximating eigenfunctions presented in [BO90] yields that
@@
$\begin{split}\norm{{u}^{(l+j)}_{H,\mathfrak{k}}-\tilde{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}&\lesssim\left(\lambda^{(l+j)}\right)^{1/2}H^{2}+\norm{\tilde{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}g(\mathfrak{k}).\end{split}$ (22)
@@
We are left to bound $\norm{\tilde{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}$. For $\mathfrak{k}$ large enough, by the coercivity of $\widetilde{a}(\cdot,\cdot)$ shown in [EGH15], for any $v\in\mathcal{H}^{1}_{\#}(\Omega)$
@@
$\begin{split}\norm{\tilde{u}_{H,\mathfrak{k}}^{(l+j)}}_{\mathcal{A}}^{2}\lesssim\;\absolutevalue{\widetilde{a}_{\mathfrak{k}}\left(\tilde{u}_{H,\mathfrak{k}}^{(l+j)},\tilde{u}_{H,\mathfrak{k}}^{(l+j)}\right)}&=\;\absolutevalue{\tilde{\lambda}_{H,\mathfrak{k}}^{(l+j)}\widetilde{m}_{\mathfrak{k}}\left(\tilde{u}_{H,\mathfrak{k}}^{(l+j)},\tilde{u}_{H,\mathfrak{k}}^{(l+j)}\right)}\\ &\leq\widetilde{C}_{\lambda}\absolutevalue{\lambda^{(l+j)}}\absolutevalue{m\left(\tilde{u}_{H,\mathfrak{k}}^{(l+j)},\tilde{u}_{H,\mathfrak{k}}^{(l+j)}\right)}=\widetilde{C}_{\lambda}\absolutevalue{\lambda^{(l+j)}}\norm{\tilde{u}_{H,\mathfrak{k}}^{(l+j)}}^{2}_{L^{2}(\Omega)}.\end{split}$ (23)
The third step follows from the bound |$\absolutevalue{\tilde{\lambda}^{(l+j)}_{H,\mathfrak{k}}}\leq c\absolutevalue{\lambda^{(l+j)}},$ for some $c\in\mathbb{R}$ by Babuška-Osborn theory. This is justified by Lemma 3.1 and the known analysis of the source problem (see above), which in particular yield the convergence $\tilde{\lambda}^{(l+j)}_{H,\mathfrak{k}}$ to $\lambda^{(l+j)}$ as $H\to 0$ and $\mathfrak{k}\to\infty$. Since $\norm{\tilde{u}_{H,\mathfrak{k}}^{(l+j)}}_{L^{2}(\Omega)}=\norm{I_{H}\tilde{u}_{H,\mathfrak{k}}^{(l+j),\text{ms}}}_{L^{2}(\Omega)}\leq\widetilde{C}\norm{\tilde{u}_{H,\mathfrak{k}}^{(l+j),\text{ms}}}_{L^{2}(\Omega)}=\widetilde{C}$ by definition, $\norm{\tilde{u}_{H,\mathfrak{k}}^{(l+j)}}_{\mathcal{A}}$ is clearly bounded above by a positive constant. Therefore, with an underlying constant that depends on the continuous eigenvalue $\lambda^{(l)}$, it holds that
$\begin{split}\norm{\tilde{u}^{(l+j)}-\tilde{u}_{H,\mathfrak{k}}^{(l+j),\rm{ms}}}_{\mathcal{A}}&\lesssim 2H^{2}+2g(\mathfrak{k})+H^{2}g(\mathfrak{k})+g^{2}(\mathfrak{k})+\left(g(\mathfrak{k})\right)^{1/2}H+H^{2}.\end{split}$